Week of 201123

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • remote: Kate (chair, DB), Vincent (security), Xavier (KIT), Alberto (monitoring), Gavin (computing), Andrew (TRIUMF), Onno (NL-T1), Christoph (CMS), Darren (RAL), Concenzio (LHCb), Ivan (ATLAS), DaveM (FNAL)

Experiments round table:

  • CMS reports ( raw view) -
    • Migration to Rucio progressing according to schedule
      • Phedex agents at sites should be switched off - Followed in a number of GGUS tickets
    • Otherwise good CPU utilization and no major troubles

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • EGI: nc
  • IN2P3:
    • CMS: PhEDEx agents have been disabled on all our VOBOXes, which will be removed this week.
    • LHCb: name space reordering tomorrow morning.
  • KISTI:nc
  • KIT:
    • Updated dCache today to 5.2.35 in specific places to deal with SVG-2020-16939 ( GOC-DB:29775).
    • On Tuesday we had a downtime for enabling TPC over xrootd for LHCb. However, as it turned out, we had to adapt the dCache configuration and restart all pools on Friday again, because what was documented for this use-case seems inaccurate.
  • NDGF: Last weeks short outage was caused by operator error. OoD set all headnodes to drain at the same time. We are looking into putting safeguards in place. (Might not be able to attend.)
  • NL-T1: NTR
  • NRC-KI: nc
  • OSG: nc
  • PIC:nc
  • RAL: NTR
  • TRIUMF: DNS server issue on Friday Nov 20, around 18:00pm (UTC) and restored at 24:00am( UTC), resulting in an unscheduled outage (https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=29787).

  • CERN computing services:
    • The final two HTCondor CEs will be upgraded to CC7 this week (ce513/514).
    • SLC6 end of life reminder - end of November - OTG:0054345.
  • CERN storage services:nc
  • CERN databases: Downtime of all integration and test Oracle databases for up to 4.5h due to HW reorganisation on Monday, Nov 30th (OTG:0060269)
  • GGUS:
    • A new release is planned for Wed this week
      • Release notes
      • A downtime has been scheduled for 07:30-09:30 UTC
      • Test alarms will be submitted as usual
  • Monitoring:
    • Distributed final SiteMon availability/reliability reports for October 2020
  • MW Officer: nc
  • Networks:nc
  • Security: NTR


Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2020-11-23 - NikolayVoytishin
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback